# Excited State Densities from Time-Dependent Density Functional Response Theory

### v1

@@ 2.1 Small-Matrix Approximation
While Eqs. 1516 provide the exact prescription for the excited-state density difference, we will make a diagonal approximation to the TDDFT matrix Eq. (3) to simplify the calculation in our examples. Neglecting the mixing between different single KS excitations is known as the Small-Matrix Approximation (SMA) 19, and taking the diagonal element of Eq. 3 yields
@@
Under this approximation, with normalization condition Eq. (12), Eq. (15) reduces to
@@
Inserting the diagonal elements of Eq. (16) directly into Eq. (20) leads to the SMA density-difference:
$\Delta n_{I}^{\mathrm{SMA}}(\mathbf{r})$ $=\frac{{G_{I}^{2}}}{\omega_{I}}\int d^{3}x\left\{\left[\left(\nu_{q}+2f_{\mathrm{Hxc}qq}(\omega_{I})\right)\Delta n_{I}^{\mathrm{KS}}(\mathbf{x})\right.+4\sum_{p\neq a}^{\infty}\frac{\nu_{q}}{\epsilon_{a}-\epsilon_{p}}{f_{\mathrm{Hxc}\,ip,ia}(\omega_{I})}{\Phi_{pa}\left(\mathbf{x}\right)}\right.$ (22)
$\left.-4\sum_{p\neq i}^{\infty}\frac{\nu_{q}}{\epsilon_{p}-\epsilon_{i}}{f_{\mathrm{Hxc}\,pa,ia}(\omega_{I})}{\Phi_{ip}\left(\mathbf{x}\right)}\right]\left(\mathbbm{1}-f_{\mathrm{Hxc}}{\chi_{s}}\right)^{-1}\left(\mathbf{x},\mathbf{r}\right)\left.+2\nu_{q}{\tilde{g}_{\mathrm{xc}\,ia,ia}\left[n_{0}\right]\left(\mathbf{x},\omega_{I}\right)}{\chi\left(\mathbf{x},r\right)}\right\}.$
@@
We note that no other approximations are applied in derivation of the Eq. (25). For an adiabatic approximation to the kernel, this result is formally equivalent to the density-difference matrix representation given by Ref. 5 in its single-transition limit; see Supplementary Material for the details.
@@ 2.1.1 Numerical Evaluation of $\Delta n^{\rm SMA}_{I}({\bf r})$
The calculation of the density difference through Eq. (16) or its SMA approximation through Eq. (22) requires addressing several issues, introducing approximations along the way.
First, we will truncate the infinite sums in Eq. (22) over all KS orbitals ($p$). Since we will be focusing only on the lower excitations, this is not a severe approximation, and convergence can be easily checked.
Second, concerning the $f_{\scriptscriptstyle\rm HXC}$ matrix elements: As noted earlier, the required matrix elements are not only between the pairs of occupied-unoccupied transitions that are the usual ingredients of linear response calculations. While the larger span does not pose a problem for adiabatic $f_{\scriptscriptstyle\rm HXC}$ approximations, computed from the second functional derivative of an energy functional, we note that these matrix elements are largely not known for the existing frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ approximations 21, 27, 23, 24, although we note that the TDDFT kernel derived from the Bethe-Salpeter equation 28, 29, 30 does contain a spatial-dependence that would provide this larger class of matrix elements. Thus when we compute the densities of excited states of double-excitation character, we will instead work directly with the excitation energies provided by the dressed frequency-dependent kernel of Refs. 21, 27, 23, 24, taking their functional derivative directly; we defer a discussion of this to Section 2.2. All other calculations will utilize an adiabatic approximation for both $f_{\scriptscriptstyle\rm HXC}(\omega)$ and $\tilde{g}_{\scriptscriptstyle\rm XC}(\omega)$. This means that the eigenvector squared value $G_{I}^{2}$ in Eq. (22) reduces to $G_{I}^{2}=1$.
@@
$\chi({\bf x},{\bf r})\approx\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})+\iint d^{3}x{{}^{\prime}}d^{3}x^{\prime\prime}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf x}{{}^{\prime}})f_{\scriptscriptstyle\rm HXC}({\bf x}{{}^{\prime}},{\bf x}^{\prime\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime\prime},{\bf r})$ (27)
@@ 2.2 Densities of States of Double-Excitation Character
With the adiabatic approximation, TDDFT is only able to predict excitations of single excitation character, composed of linear combinations of one electron promoted from an occupied KS state to a virtual one. The formulation for excited-state densities in Refs. 4, 5 is restricted to this case. For states of double-excitation character, a frequency-dependent kernel was developed 20, 21, 22, 23, 24 that has been shown to perform well on a range of systems 32, most recently capturing the curve-crossing between the 1Bu and 2Ag singlet states in trans-butadiene 24, for example. The kernel of Ref. 21, 27 operates within a dressed single-pole approximation (DSPA), generalized to a dressed Tamm-Dancoff approximation (DTDA) when more than one single excitation couples to a double, while, in order to obtain meaningful oscillator strengths as well as energies, the kernel of Ref. 23, 24 operates within a dressed small-matrix approximation (DSMA), generalized to the fully dressed TDDFT (DTDDFT). While Eq. (15) and its SMA Eq (22) hold also beyond the adiabatic approximation, they require matrix elements of the kernel that in a larger space than those derived in the approximations, as discussed above. Therefore, we instead take the functional derivative of Eq. 8 in a different way, directly from the DSPA and DSMA expressions for the excitation energy.
We begin with the DSMA case where 23
@@
Eq. (29) is given for the “0” variant of DSMA Hxc kernel from the work23; other variants reduce the number of two-electron integrals needed by replacing some of the Hamiltonian matrix elements with either Kohn-Sham or appropriate adiabatic TDDFT values. For example, in the “A” variant, $H_{qq}-H_{00}\to\omega_{q}^{\mathrm{ASMA}}$ and $H_{DD}-H_{00}\to\omega_{jb}^{\mathrm{ASMA}}+\omega_{kc}^{\mathrm{ASMA}}$. Here, we have chosen to use the first replacement, but we keep the Hamiltonian matrix element for the double-excitation part. But we note that the variants all give similar results for this system. The two excitation energies given by DSMA Eq. (28) are therefore
@@
Taking the derivative Eq. 8 of Eq.(30) then leads to the DSMA excited state density:
$\Delta n^{\mathrm{DSMA}}_{\pm}({\bf r})$ $=\frac{1}{2}(1\pm\cos\theta)\Delta n^{\mathrm{ASMA}}({\bf r})$ (31)
$+\frac{1}{2\Omega_{\pm}}\left\{\frac{1}{2}(1\mp\cos\theta)\frac{\delta\left(H_{DD}-H_{00}\right)^{2}}{\delta v_{\mathrm{ext}}({\bf r})}\right.$
$\left.+\frac{\delta H_{qD}^{2}}{\delta v_{\mathrm{ext}}({\bf r})}\pm\sin\theta\frac{\delta}{\delta v_{\mathrm{ext}}({\bf r})}\left[H_{qD}\left(\omega^{\mathrm{ASMA}}+(H_{DD}-H_{00})\right)\right]\right\}.$
In the first term above the adiabatic excited state density difference $\Delta n^{\mathrm{ASMA}}$ appears, which we obtain from Eq. (22), with the kernel $f_{\scriptscriptstyle\rm HXC}(\omega_{I})$ approximated by the chosen adiabatic approximation. This density contributes through the mixing angle $\theta$ defined via
@@
Note that in the first term of Eq. (31), the adiabatic density-difference $\Delta n^{\mathrm{ASMA}}({\bf r})$ is scaled by the square of the Casida eigenvector $G_{\pm}^{2}$ from Eq. (21),
@@
which is a measure of the amount of single-excitation character of the state 23. In the limit $\cos\theta=1$, the state indexed by $+$ reduces to a pure double excitation while that indexed by -- reduces to a pure single excitation; the opposite is true for $\cos\theta=-1$. On the other hand, when $\cos\theta\to 0$, the states gain an equal contribution of single and double excitation character, and become nearly degenerate. Their densities however remain distinct, due to the last term in Eq. (31).
For the second and third terms of Eq. (31), which contain non-adiabatic contributions to the density difference, one needs derivatives of the Hamiltonian matrix elements. These can be evaluated from evaluating the Hamiltonian matrix elements via Slater-Condon rules for one-body and two-body integrals and then taking the derivatives (details in Supplementary Material):
@@
While including both excitation and de-excitation components to the TDDFT matrix is necessary to obtain meaningful transition densities 23, we will check whether including only the excitation components is adequate for the excited state density itself, by using the Tamm-Dancoff approximation of DSMA. This is just the dressed single-pole approximation (DSPA) that was derived earlier in Ref. 21:
@@
The excited-state density difference following from Eq. (36) is
@@
±$\pm\sin\theta\frac{\delta H_{qD}}{\delta v_{\rm ext}({\bf r})},$
@@ 3 Results and Analysis on Model Systems
We now apply our method to a set of one-dimensional models of two soft-Coulomb interacting electrons with different external potentials. We define the three following external potentials: (i) Helium $v_{\rm ext}^{\mathrm{1DHe}}(x)=\frac{-2}{\sqrt{1+x^{2}}}$ to model local excitations, (ii) two double-well models, one in which the ground-state density is largely localized in a soft-Coulomb potential on the left, $v_{\rm ext}^{\mathrm{gs-soft}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{1}{\cosh^{2}(x-R/2)}$ 33, and the other in which the ground-state density is more tightly localized in the left well, $v_{\rm ext}^{\mathrm{gs-loc}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{2.9}{\cosh^{2}(x+R/2)}-\frac{1}{\cosh^{2}(x-R/2)}$, where $R=7$a.u., and in these wells we will model charge-transfer excitations, and (iii) perturbed harmonic $v_{\rm ext}^{\mathrm{Harm_{\gamma}}}(x)=\frac{1}{2}x^{2}+\gamma|x|$ for modeling the double excitations, where $\gamma$ parameter tunes the degree of mixing between a single excitation and double-excitation in the second multiplet.
@@
With just two electrons, the Kohn-Sham ground-state has both electrons occupying the lowest-energy orbital $i=1$, and the KS response function for all three cases can be written as
$\chi_{\scriptscriptstyle\rm S}(x,x{{}^{\prime}})=-4\sum_{a}^{\mathrm{unocc}}\frac{\phi_{i}(x)\phi_{a}(x)\phi_{a}(x{{}^{\prime}})\phi_{i}(x{{}^{\prime}})}{\epsilon_{a}-\epsilon_{i}}.$ (40)
@@
The simplicity of the models enables the computations of the exact KS orbitals, and we can compare the effect of using approximate versus exact KS orbitals in the expressions for the excited-state densities. We will first consider the lowest energy excitations in 1DHe and the two double-wells. These are dominated by single excitations, and so we will consider two adiabatic approximations: exact-exchange (EXX) and local-density approximation (LDA) 34, which are first-principles approximations at two opposite extremes, the former derived from many-body perturbation theory has non-local density-dependence and the latter derived from the uniform electron gas paradigm has completely local density-dependence. We will compare the fully-converged SMA densities (Eq. 22) with those obtained from STL (Eq. 25). For the Harmγ system, we compare the performance of ASMA using the EXX approximation with the DSMA Eq. (31).
@@
For our calculations of the exact ground and excited-state densities, and the exact, Kohn-Sham, EXX, and LDA orbitals of these one-dimensional systems, we use the Octopus code35. The exact Kohn-Sham potential and orbitals are obtained from first inverting the ground-state Kohn-Sham equation using the exact ground-state density of a two-electron system, and then finding its eigenstates and eigenvalues in an in-house code. The helium model was simulated in a box ranging from -40 to 40 a.u. with a grid spacing of 0.1 a.u. For both double-well models, a box from -50 to 50 a.u. with the same grid spacing was used. Finally, for the system in the harmonic potential, simulations were performed in a box from -20 to 20 a.u. with a grid spacing of 0.05 a.u.

### v2

@@ 2.1 Small-Matrix Approximation
While Eqs. 121516 provide the exact prescription for the excited-state density difference, we will make a diagonal approximation to the TDDFT matrix Eq. 3 to simplify the calculation in our examples. Neglecting the mixing between different single KS excitations is known as the Small-Matrix Approximation (SMA) 29, and taking the diagonal element of Eq. 3 yields
@@
The utilization of SMA for the excited-state densities has limitations similar to those observed for the excitation frequencies: it performs best when the excited state is dominated by a single KS transition. If the excitation involves a significant contribution from multiple KS states, the densities arising from SMA will be less accurate.
Under SMA, with normalization condition Eq. 12, Eq. 15 reduces to
@@
Inserting the diagonal elements of Eq. 16 directly into Eq. 20 leads to the SMA density-difference:
$\Delta n_{I}^{\mathrm{SMA}}(\mathbf{r})$ $=\frac{{G_{I}^{2}}}{\omega_{I}}\int d^{3}x\left\{\left[\left(\nu_{q}+2{f_{\scriptscriptstyle\rm HXC}}_{qq}(\omega_{I})\right)\Delta n_{I}^{\mathrm{KS}}(\mathbf{x})\right.+4\sum_{p\neq a}^{\infty}\frac{\nu_{q}}{\epsilon_{a}-\epsilon_{p}}{{f_{\scriptscriptstyle\rm HXC}}_{ip,ia}(\omega_{I})}{\Phi_{pa}\left(\mathbf{x}\right)}\right$. (22)
$\left.-4\sum_{p\neq i}^{\infty}\frac{\nu_{q}}{\epsilon_{p}-\epsilon_{i}}{{f_{\scriptscriptstyle\rm HXC}}_{pa,ia}(\omega_{I})}{\Phi_{ip}\left(\mathbf{x}\right)}\right]\left(\mathbbm{1}-{f_{\scriptscriptstyle\rm HXC}}{\chi_{s}}\right)^{-1}\left(\mathbf{x},\mathbf{r}\right)\left.+2\nu_{q}{\tilde{g}_{\mathrm{xc},\,ia,ia}\left[n_{0}\right]\left(\mathbf{x},\omega_{I}\right)}{\chi\left(\mathbf{x},r\right)}\right\}.$
@@
We note that no other approximations are applied in derivation of the Eq. 25. For an adiabatic approximation to the kernel, this result is formally equivalent to the density-difference matrix representation given by Ref. 5 in its single-transition limit; see Supplementary Material for the details.
@@ 2.1.1 Numerical Evaluation of $\Delta n^{\rm SMA}_{I}({\bf r})$
The calculation of the density difference through Eq. 16 or its SMA approximation through Eq. 22 requires addressing several issues, introducing approximations along the way.
First, we will truncate the infinite sums in Eq. 22 over all KS orbitals ($p$). Since we will be focusing only on the lower excitations, this is not a severe approximation, and convergence can be easily checked.
Second, concerning the $f_{\scriptscriptstyle\rm HXC}$ matrix elements: As noted earlier, the required matrix elements are not only between the pairs of occupied-unoccupied transitions that are the usual ingredients of linear response calculations. The larger span does not pose a problem for adiabatic $f_{\scriptscriptstyle\rm HXC}$ approximations: because this is obtained from the second functional derivative of a ground-state energy functional, its form in real-space is known so any matrix element could be computed. However, it is a problem for existing frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ approximations, such as those relevant to states of double-excitation character, because only the matrix elements involving the KS states that significantly contribute to these states are known 31, 37, 33, 34. We note that the TDDFT kernels derived from the Bethe-Salpeter equation (BSE) 38, 39, 40 do provide this larger class of matrix elements, however, their practical use is limited due to computational cost. Unlike most known TDDFT kernels, the frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ arising from contracting the four-point BSE kernel does not explicitly depend on the electron density, and functional derivatives with respect to the external potential of the resulting matrix elements would be directly computed.
Thus when we compute the densities of excited states of double-excitation character, we will instead work directly with the excitation energies provided by the dressed frequency-dependent kernel of Refs. 31, 37, 33, 34, taking their functional derivative directly; we defer a discussion of this to Section 2.2. All other calculations will utilize an adiabatic approximation for both $f_{\scriptscriptstyle\rm HXC}(\omega)$ and $\tilde{g}_{\scriptscriptstyle\rm XC}(\omega)$. This means that the eigenvector squared value $G_{I}^{2}$ in Eq. 22 reduces to $G_{I}^{2}=1$.
@@
$\chi({\bf x},{\bf r})\approx\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})+\iint d^{3}x{{}^{\prime}}d^{3}x^{\prime\prime}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf x}{{}^{\prime}})f_{\scriptscriptstyle\rm HXC}({\bf x}{{}^{\prime}},{\bf x}^{\prime\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime\prime},{\bf r}).$ (27)
In most of the cases we have studied, the terms of higher order in $f_{\scriptscriptstyle\rm HXC}$ of the series Eq. 2627 bring a negligibly small correction to the density difference, and we expect this to be generally true for weakly interacting systems. However, as we show in Sec. 3.2, an inadequate KS approximation of the ground state quantities of a system can lead to the divergence of the first order Taylor expansion through the KS linear response function $\chi_{\scriptscriptstyle\rm S}$ if the KS gaps are significantly underestimated.
@@ 2.2 Densities of States of Double-Excitation Character
With the adiabatic approximation, TDDFT is only able to predict excitations of single excitation character, composed of linear combinations of one electron promoted from an occupied KS state to a virtual one. The formulation for excited-state densities in Refs. 4, 5 is restricted to this case. For states of double-excitation character, a frequency-dependent kernel was developed 30, 31, 32, 33, 34 that has been shown to perform well on a range of systems 42, most recently capturing the curve-crossing between the 1Bu and 2Ag singlet states in trans-butadiene 34, for example. The kernel of Ref. 31, 37 operates within a dressed single-pole approximation (DSPA), generalized to a dressed Tamm-Dancoff approximation (DTDA) when more than one single excitation couples to a double, while, in order to obtain meaningful oscillator strengths as well as energies, the kernel of Ref. 33, 34 operates within a dressed small-matrix approximation (DSMA), generalized to the fully dressed TDDFT (DTDDFT). While Eq. 15 and its SMA Eq (22) hold also beyond the adiabatic approximation, they require matrix elements of the kernel that in a larger space than those derived in the approximations, as discussed above. Therefore, we instead take the functional derivative of Eq. 8 in a different way, directly from the DSPA and DSMA expressions for the excitation energy.
We begin with the DSMA case where 33
@@
Eq. 29 is given for the “0” variant of DSMA HXC kernel from the work33; other variants reduce the number of two-electron integrals needed by replacing some of the Hamiltonian matrix elements with either Kohn-Sham or appropriate adiabatic TDDFT values. For example, in the “A” variant, $H_{qq}-H_{00}\to\omega_{q}^{\mathrm{ASMA}}$ and $H_{DD}-H_{00}\to\omega_{jb}^{\mathrm{ASMA}}+\omega_{kc}^{\mathrm{ASMA}}$. Here, we have chosen to use the first replacement, but we keep the Hamiltonian matrix element for the double-excitation part. But we note that the variants all give similar results for the system we will study. The two excitation energies given by DSMA Eq. 28 are therefore
@@
Taking the derivative Eq. 8 of Eq. 30 then leads to the DSMA excited state density:
$\Delta n^{\mathrm{DSMA}}_{\pm}({\bf r})$ $=\frac{1}{2}(1\pm\cos\theta)\frac{\omega^{\mathrm{ASMA}}}{\Omega_{\pm}}\Delta n^{\mathrm{ASMA}}({\bf r})$ (31)
$+\frac{1}{2}(1\mp\cos\theta)\frac{H_{DD}-H_{00}}{\Omega_{\pm}}\frac{\delta\left(H_{DD}-H_{00}\right)}{\delta v_{\mathrm{ext}}({\bf r})}$
$+\frac{1}{2\Omega_{\pm}}\left\{\frac{\delta H_{qD}^{2}}{\delta v_{\mathrm{ext}}({\bf r})}\pm\sin\theta\frac{\delta}{\delta v_{\mathrm{ext}}({\bf r})}\left[H_{qD}\left(\omega^{\mathrm{ASMA}}+(H_{DD}-H_{00})\right)\right]\right\}.$
In the first term above the adiabatic excited state density difference $\Delta n^{\mathrm{ASMA}}$ appears, which we obtain from Eq. 22, with the kernel $f_{\scriptscriptstyle\rm HXC}(\omega_{I})$ approximated by the chosen adiabatic approximation. This density contributes through the mixing angle $\theta$ defined via
@@
Note that in the first term of Eq. 31, the adiabatic density-difference $\Delta n^{\mathrm{ASMA}}({\bf r})$ is scaled by the square of the Casida eigenvector $G_{\pm}^{2}$ from Eq. 21,
@@
which is a measure of the amount of single-excitation character of the state 33. In the limit $\cos\theta=1$, the state indexed by ++ reduces to a pure double excitation while that indexed by - reduces to a pure single excitation; the opposite is true for $\cos\theta=-1$. On the other hand, when $\cos\theta\to 0$, the states gain an equal contribution of single and double excitation character, and become nearly degenerate. Their densities however remain distinct, due to the last term in Eq. 31.
For the second and third terms of Eq. 31, which contain non-adiabatic contributions to the density difference, one needs derivatives of the Hamiltonian matrix elements. These can be evaluated from evaluating the Hamiltonian matrix elements via Slater-Condon rules for one-body and two-body integrals and then taking the derivatives (details in Supplementary Material):
@@
While including both excitation and de-excitation components to the TDDFT matrix is necessary to obtain meaningful transition densities 33, we will check whether including only the excitation components is adequate for the excited state density itself, by using the Tamm-Dancoff approximation of DSMA. This is just the dressed single-pole approximation (DSPA) that was derived earlier in Ref. 31:
@@
The excited-state density difference following from Eq. 36 is
@@
$\pm\sin\theta\frac{\delta H_{qD}}{\delta v_{\rm ext}({\bf r})},$
@@
It is important to note that, although the factors $\frac{1}{2}(1\pm\cos\theta)$ appear in the DSPA density difference Eq. 38, they are not the same as $G_{I}^{2}$ of DSPA, since Eq. 33 is derived from the normalization Eq. 12 when the DSMA kernel is used.
@@ 3 Results and Analysis on Model Systems
We now apply our method to a set of one-dimensional models of two soft-Coulomb interacting electrons with different external potentials. We define the three following external potentials: (i) 1DHe: one-dimensional Helium $v_{\rm ext}^{\mathrm{1DHe}}(x)=\frac{-2}{\sqrt{1+x^{2}}}$ to model local excitations, (ii) two double-well models, one in which the ground-state density is largely localized in a soft-Coulomb potential on the left, $v_{\rm ext}^{\mathrm{gs-soft}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{1}{\cosh^{2}(x-R/2)}$ 43, and the other in which the ground-state density is more tightly localized in the left well, $v_{\rm ext}^{\mathrm{gs-loc}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{2.9}{\cosh^{2}(x+R/2)}-\frac{1}{\cosh^{2}(x-R/2)}$, where R=7R=7a.u., and in these wells we will model charge-transfer excitations, and (iii) Harmγ: perturbed harmonic $v_{\rm ext}^{\mathrm{Harm_{\gamma}}}(x)=\frac{1}{2}x^{2}+\gamma|x|$ for modeling the double excitations, where $\gamma$ parameter tunes the degree of mixing between a single excitation and double-excitation in the second multiplet.
@@
With just two electrons, the Kohn-Sham ground-state has both electrons occupying the lowest-energy orbital i=0{i=0}, and the KS response function for all three cases can be written as
$\chi_{\scriptscriptstyle\rm S}(x,x{{}^{\prime}})=-4\sum_{a}^{\mathrm{unocc}}\frac{\phi_{0}(x)\phi_{a}(x)\phi_{a}(x{{}^{\prime}})\phi_{0}(x{{}^{\prime}})}{\epsilon_{a}-\epsilon_{0}}.$ (40)
@@
The simplicity of the models enables the computations of the exact KS orbitals, and we can compare the effect of using approximate versus exact KS orbitals in the expressions for the excited-state densities. We will first consider the lowest energy excitations in 1DHe and the two double-wells. These are dominated by single excitations, and so we will consider two adiabatic approximations: exact-exchange (EXX) and local-density approximation (LDA) 44, which are first-principles approximations at two opposite extremes, the former derived from many-body perturbation theory has non-local density-dependence and the latter derived from the uniform electron gas paradigm has completely local density-dependence. We will compare the fully-converged SMA densities (Eq. 22) with those obtained from STL (Eq. 25). For the Harmγ system, we compare the performance of ASMA using both the EXX and LDA approximations with the DSMA Eq. 31.
@@
For our calculations of the exact ground and excited-state densities, and the exact, Kohn-Sham, EXX, and LDA orbitals of these one-dimensional systems, we use the Octopus code45. The exact Kohn-Sham potential and orbitals are obtained from first inverting the ground-state Kohn-Sham equation using the exact ground-state density of a two-electron system, and then finding its eigenstates and eigenvalues in an in-house code. The 1DHe model was simulated in a box ranging from -40 to 40 a.u. with a grid spacing of 0.1 a.u. For both double-well models, a box from -50 to 50 a.u. with the same grid spacing was used. Finally, for the system in the harmonic potential, simulations were performed in a box from -20 to 20 a.u. with a grid spacing of 0.05 a.u.